Can someone help with this problem?
If a line passes through (2,6) and has the same y-intercept as a line whose equation is x-3y=18
and
If a line passes through (2,4) and has the same y-intercept as a line whose equation is x-4y=8.
I need the answer is slope intercept form

Answer by stanbon(48558)

1.If a line passes through (2,6) and has the same y-intercept as a line whose equation is x-3y=18
---------
The given line has y-intercept (0,-6)
The equation you want is y=mx+b and you need to find m and b:
m=[6--6]/[2-0]=6
You have y=6,x=2,m=6 so you can solve for b as follows:
6=6(2)+b; b=-6
EQUATION: y=6x-6
and
2. If a line passes through (2,4) and has the same y-intercept as a line whose equation is x-4y=8.
------------
Rewrite that equation as 4y=x-8; y=(1/4)x-2
The given line has slope (1/4)
You have y=4,x=2 and m=(1/4); solve for b as follows:
4=(1/4)2+b
4=(1/2)+b
b=7/2
EQUATION:
y=(1/4)x+7/2
Cheers,
Stan H.
Cheers,
Stan H.